# Unveiling User Preferences: A Knowledge Graph and LLM-Driven Approach for Conversational Recommendation

### v1

@@ 4.2.2. Comparison of Enhancement Methods
Figure 3. Comparison of recommendation performance on the INSPIRED dataset with different enhancers. Green percentages show improvements over baselines.
[…]
[Performance comparison of recommendation models]The figure shows two bar charts comparing HR@10 and HR@50 for four recommendation models (KBRD, KGSF, BERT, Llama3.1-8B) and their enhancements. Each model has four bars: Base, Llama-Summary, GPT-4o, and COMPASS. COMPASS consistently outperforms other methods across all models, with the most significant improvements for KBRD and KGSF. Percentage improvements are displayed above each enhanced bar.
Table 2. Ablation study on recommendation task performance. The best-performing results are highlighted in bold.
@@
To evaluate COMPASS’s effectiveness, we compare it with Llama-Summary and GPT-4o as baseline enhancers, which generate user preference summaries without KG access. Figure 3 illustrates the performance of these methods across various CRS models on the INSPIRED dataset. While all enhancement methods show improvements over the base models, COMPASS consistently outperforms both alternatives. Notably, COMPASS consistently outperforms GPT-4o across all models, with improvements ranging from marginal (for Llama3.1-8B) to substantial (for KBRD and KGSF). This superior performance, achieved despite GPT-4o being a much larger model, demonstrates the importance of integrating domain-specific knowledge into preference modeling. The substantial performance gap between COMPASS and Llama-Summary highlights the effectiveness of our framework in leveraging both structured knowledge and LLM capabilities. These results show COMPASS’s ability to generate more accurate and contextually relevant user preference summaries, leading to improved recommendation performance across different CRS architectures.
4.2.3. Ablation study
The results in Table 2 show that COMPASS generally achieves the best performance across most metrics, highlighting its superior capability. The COM w/o REC variant shows the largest performance drop, emphasizing the crucial role of including recommended items in preference summaries to guide downstream models. Both COM w/o KG and COM w/o GEP variants demonstrate comparable performance declines, highlighting the crucial roles of KG integration for understanding item relationships and GEP for comprehending KG-augmented context. These findings show the importance of each COMPASS component, with their synergy driving the superior performance of the full model.
4.3. Preference Generation Evaluation
This study introduces a novel approach to enhance CRS by generating user preference summaries. As the first attempt in this direction, this evaluation focuses on comparing COMPASS with its variants to assess the effectiveness of its key components. Additionally, a case study that includes examples from Llama-Summary and GPT-4o provides additional qualitative insights.
4.3.1. Automatic Evaluation
Table 3. Evaluation of generated user preference summaries. RP is reasoning proficiency and FC is factual consistency. The best results are highlighted in bold.
@@ 4.3.1. Automatic Evaluation
Table 3 presents a comparative analysis of COMPASS and its variants, COM w/o KG and COM w/o GEP, on the ReDial and INSPIRED datasets. COMPASS consistently outperforms its variants across all metrics on both datasets. Lexical similarity metrics (ROUGE scores) demonstrate that COMPASS generates summaries that align closely with reference texts, while higher reasoning proficiency and factual consistency scores illustrate its enhanced quality of reasoning and factual accuracy.
The performance of COMPASS compared to COM w/o KG indicates the value of KG integration in enhancing summary quality. This suggests that incorporating structured knowledge grounds the model’s output in domain-specific information, which is crucial for accurate preference generation. Moreover, the comparative performance of COM w/o GEP and COMPASS indicates that, despite the presence of structured knowledge from the KG, there remains a modality gap between graph structures and natural language, posing challenges for LLMs in fully understanding and utilizing this information. These results validate the COMPASS framework, and demonstrating the synergistic effect of combining KGs with LLMs, and showing the critical role of our pre-training strategy in enhancing user preference generation.
4.3.2. LLM-Simulated User Evaluation
(a) ReDial: Alignment
(b) ReDial: Explainability
(c) INSPIRED: Alignment
(d) INSPIRED: Explainability
Figure 4. LLM-simulated user rankings for User Preference Alignment and Explainability across ReDial and INSPIRED datasets.
To assess user-perceived quality, we employ LLM-simulated user evaluations on explainability and user preference alignment. Figure 4 presents the results across the ReDial and INSPIRED datasets. COMPASS is consistently preferred over COM w/o KG and COM w/o GEP in both metrics. On the ReDial dataset, COMPASS is most preferred in in approximately 75% of the cases for both alignment and explainability. Although the preference gap is smaller on INSPIRED, COMPASS maintains a clear advantage. Notably, COM w/o NP is consistently the least preferred, showing the importance of pre-training in generating meaningful preference summaries. These results demonstrate COMPASS’s superior ability to generate preference summaries that align closely with user preferences while providing clear, user-friendly explanations for recommendations, enhancing both fidelity and transparency in CRSs.
4.3.3. Case Study
We present a detailed case study in Table 5 to visualize a sample dialogue along with the corresponding preference summaries generated by different models. It demonstrates that COMPASS accurately identifies the key information for capturing user preference and provides interpretable summaries. The detailed analysis can be found in Appendix C.2.
5. Conclusion
In this paper, we introduce COMPASS, a novel framework that synergizes KGs and LLMs to unveil user preferences in Conversational Recommender Systems (CRSs). To address the modality gap between structured knowledge and natural language, we propose a graph entity captioning mechanism that transforms KG structures into LLM-compatible representations. Through knowledge-aware instruction tuning, COMPASS becomes proficient in performing cross-modal reasoning, generating comprehensive and interpretable user preference summaries. Our proposed framework has been extensively evaluated as a plug-and-play enhancement for various existing CRS models across benchmark datasets. The results demonstrate COMPASS’s effectiveness in significantly improving both the recommendation performance and the explainability of these base models. The adaptive integration mechanism of COMPASS allows for seamless enhancement of diverse CRS architectures without requiring structural modifications, showcasing its versatility and potential for widespread adoption in the field of CRSs.
Appendix A Related Work
A.1. Conversational Recommender System
Conversational Recommender Systems (CRSs) represent an evolution in recommendation technology, designed to understand user preferences through interactive dialogue and provide personalized recommendations (Jannach et al., 2021). These systems face a unique challenge in modeling user preferences from typically brief and sparse conversational data. To address this, recent works incorporate external knowledge sources into CRSs to enhance preference modeling. Knowledge graphs (KG) enable systems to infer implicit user interests and capture complex relationships between items (Chen et al., 2019; Zhou et al., 2020), while user reviews provide additional insights into item attributes and user opinions (Lu et al., 2021; Zhou et al., 2022; Kim et al., 2023), helping to address the challenge of limited contextual information. Moreover, the integration of pre-trained language models (PLMs) such as DialoGPT (Zhang et al., 2020; Wang et al., 2022c, a) and BART (Lewis et al., 2020; Qiu et al., 2024) has further improved contextual understanding and language generation capabilities in CRSs.
More recently, the advent of powerful LLMs has introduced new opportunities for CRSs. Studies have focused on improving LLM evaluation through enhanced user simulators (Wang et al., 2023b; Zhu et al., 2024), exploring zero-shot recommendation capabilities (He et al., 2023), and utilizing LLMs for task planning in CRSs (Feng et al., 2023; Fang et al., 2024; Li et al., 2024). However, one critical aspect that remains underexplored is the role of explainable user preference in CRSs. Although some conceptual designs propose using LLMs to manage user preferences (Friedman et al., 2023), they lack quantitative validation. (Xi et al., 2024) addresses user preference memories in sequential CRS but does not fully harness LLM’s reasoning abilities for preference generation. Our work is the first to introduce a framework that synergizes the LLM and the KG for unveiling user preferences in CRSs, enhancing both preference modeling and recommendation explainability.
A.2. Explainable Recommendation
Explainable recommendation has emerged as a crucial areas of research, aiming to enhance user trust, satisfaction, and decision making by providing transparent rationales for recommendations. Early research focused on leveraging user-item interactions and content features to generate simple explanations (Zhang and Chen, 2020), using techniques like recurrent neural networks (Donkers et al., 2017), attention mechanisms (Chen et al., 2018), and graph neural networks (Wang et al., 2019). Recent developments in LLMs have opened new avenues for explainable recommendation. Studies (Ma et al., 2024; Wang et al., 2024) have demonstrated LLMs’ potential in generating explanations by analyzing user’s past interactions. However, current LLM-based methods primarily operate within the constraints of static user profiles, where preferences are inferred from historical behavior. Our work introduces a novel LLM-KG framework for CRSs, enhancing both preference modeling and recommendation explainability in dynamic interaction contexts.
Appendix B Details of Experimental Setting
B.1. Model Details
B.1.1. Backbone Models
The detailed information for each backbone is as follows:
  • ReDial (Li et al., 2018): Introduces the ReDial dataset and model for conversational recommendation, utilizing a denoising auto-encoder(He et al., 2017) as the recommendation module.
  • KBRD (Chen et al., 2019): This method leverages an external KG to enhance the semantics of entities mentioned in the dialogue history.
  • KGSF (Zhou et al., 2020): This method employs a dual KG approach, integrating semantic information from both word-level and entity-level KGs. It applies a mutual information maximization technique to align these two semantic spaces, enhancing the overall representation.
  • BERT (Devlin et al., 2019): This method utilizes a bidirectional Transformer model (Vaswani et al., 2017) pre-trained on a large-scale corpus. We use the [CLS] token representation for the recommendation task.
  • GPT-2 (Radford et al., 2019): An auto-regressive language model. We fine-tune it on dialogue history and use the last token representation for the recommendation task.
  • Llama3.1-8B (Dubey et al., 2024): This is an open-source LLM that has been trained using both supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022). For the recommendation task, we fine-tune LLama3.1-8B using LoRA on the dialogue history and use the last token representation as the user preference embedding.
  • BARCOR (Wang et al., 2022b): This model introduces a unified framework that integrates BART with a knowledge graph, enabling both recommendation and response generation within a single model.
  • PECRS (Ravaut et al., 2024): This model integrates recommendation and response generation into a single training phase using LoRA (Hu et al., 2022) with a pre-trained language model, eliminating the need for separate modules or external knowledge graphs.
This study introduces a novel approach to enhance CRS by generating user preference summaries. As there are no existing enhancers specifically designed for this task, we use Llama3.1-8B [^5] https://ollama.com/ and GPT-4o666gpt-4o-2024-05-13 as baselines for comparison, both of which generate user preference summaries without access to KG information. To clarify, Llama3.1-8B is used in two different capacities in this work: first, as a backbone CRS model adapted through training for the recommendation task, and second, as Llama-Summary for generating user preference summaries without KG access.
B.1.2. COMPASS Integration
We enhance CRS models performance by incorporating our COMPASS framework as described in Section 3.4. The integration method is consistent across KBRD, KGSF, and BERT, utilizing the adaptive gating mechanism to combine COMPASS-generated preference summaries with the original model representations. For preference-enhanced Llama3, we employ a different integration strategy. We directly concatenate the generated user preference summary with the conversation history. This approach leverages Llama3.1-8B’s strong language understanding capabilities to process the additional preference information.
B.2. Instruction Design and Caption Templates
The instruction templates for graph entity captioning and knowledge-aware instruction tuning are presented in Tables 6,  7, and 8. Table 6 shows various instruction templates for different node types in the knowledge graph, including non-item nodes (actors, directors, genres, keywords, production companies) and item nodes (movies). Table 7 presents the structured template for user preference modeling and recommendation, designed to produce JSON-formatted output. Table 8 displays the instruction for Llama3.1-8B to make direct movie recommendations.
The caption templates for various entity types in the KG (such as genres, actors, directors, companies, and keywords) are presented in Table 9. These templates guide the translation of graph structures into natural language descriptions, emphasizing each entity’s relationships within the domain. The templates enable COMPASS to generate contextual descriptions that preserve graph structure while presenting information in a format compatible with the LLM, facilitating effective cross-modal analysis.
B.3. Ground Truth User Preference Summary Generation
To generate reliable ground truth user preference summaries, we employed OpenAI’s GPT-4o model777gpt-4o-2024-05-13, with temperature set to 0 for deterministic outputs. The model performed cross-modal reasoning by integrating the complete dialogue history with structured metadata of items mentioned in the conversation, derived from the KG. We developed a structured instruction template, presented in Table 10, to guide this analysis process. This template outlines a step-by-step approach for discerning user preferences through conversation analysis, KG cross-referencing, and insight synthesis, resulting in a comprehensive user preference summary.
B.4. Detailed Evaluation Process
B.4.1. Semantic Understanding Evaluation
Following the method proposed by (Wang et al., 2023a), we adopt an LLM-based evaluation approach where GPT-4o-mini serves as an automated scorer to evaluate the reasoning proficiency (RP) and fact consistency (FC) of the user preference summaries generated. The evaluation of RP is centered on three main criteria: (1) Logical Coherence, which assesses the internal consistency of the reasoning process; (2) Accuracy of Inferences, which measures how precisely deductions are made from the dialogue and KG data; and (3) Relevance of Recommendations, which evaluates whether the suggested items are appropriate given the user’s expressed interests. FC evaluation similarly considers three aspects: (1) Dialogue Alignment, which measures fidelity to user-expressed preferences; (2) KG Consistency, assessing the accuracy of preference keywords relative to the KG content; and (3) Claim Validity, checking for unsupported assertions while allowing for reasonable expansions. The evaluation uses the dialogue history, KG data related to the mentioned entities in the dialogue, and the generated preference summary as input. GPT-4o-mini then analyzes these inputs and generates scores for RP and FC on a scale of 0-100. The average scores are calculated for all test samples.
B.4.2. User-Centric Evaluation
To complement the quantitative metrics with user-oriented insights, a user-centric evaluation is performed using GPT-4o-mini as a simulated user, following the approach outlined by (Ji et al., 2023; Gao et al., 2024). We evaluate the Explainability and User Preference Alignment of the generated preference summaries and recommendations. Explainability refers to how clearly the system justifies its summaries and recommendations in a way that is easily understandable from the user’s perspective. User Preference Alignment assesses how accurately the system captures and responds to the preferences expressed by the user in the conversation. The evaluation uses the dialogue history and preference summaries from different models as inputs. GPT-4o-mini ranks these summaries according to the EX and UPA criteria, providing justifications for each ranking.
Appendix C Further Experiment
C.1. Evaluation of Preference Encoding Methods
Table 4. Comparison of preference encoding methods for integrating COMPASS generated summaries. The best results are highlighted in bold.
Model ReDial
HR@10 HR@50 NDCG@10 NDCG@50 MRR@10 MRR@50
KBRD 0.151 0.336 0.099 0.136 0.071 0.079
   +EOS 0.188 0.379 0.099 0.141 0.073 0.082
   +Text 0.199 0.412 0.103 0.150 0.075 0.085
KGSF 0.183 0.378 0.098 0.140 0.072 0.081
   +EOS 0.190 0.383 0.102 0.144 0.075 0.084
   +Text 0.198 0.413 0.105 0.152 0.076 0.088
BERT 0.143 0.319 0.073 0.108 0.052 0.059
   +EOS 0.162 0.348 0.084 0.125 0.060 0.069
   +Text 0.182 0.382 0.098 0.142 0.073 0.082
Our COMPASS framework demonstrates versatility in generating user preference summaries, offering both comprehensive textual descriptions and compact EOS token embeddings. To integrate these outputs into existing CRS models, we evaluated two encoding methods: a lightweight encoder and a PLM-based text encoder. Table 4 presents results for KBRD, KGSF, and BERT models on the ReDial dataset, which shows the effectiveness of both approaches. The PLM-based text encoder (108M parameters) excels in processing COMPASS’s textual summaries, achieving superior performance across all metrics.
Conversely, the lightweight encoder (16.8M parameters, $\sim$15.5of PLM-based) efficiently processes COMPASS’s EOS token embeddings. Despite its simplicity, it significantly improves on the baseline performance, increasing KBRD’s HR@10 and HR@50 by 24.5% and 12.8% respectively. This demonstrates COMPASS’s capability to distill user preferences into compact yet informative representations. Both methods enhance performance across different base models, highlighting COMPASS’s adaptability in generating preference summaries. This versatility enables flexible deployment across various computational environments, from resource-constrained systems to high-performance platforms, thereby accommodating a wide range of application scenarios.
Table 5. Case Study: Comparison of preference summaries generated by different models. Green shows information supported by the knowledge graph, yellow for correct inferences that are not explicitly detailed in the knowledge graph but align with the content, and red for unsupported and unreasonable details
@@ C.1. Evaluation of Preference Encoding Methods
COM w/o KG: Reasoning: The user mentioned ‘The Professional (1981)’ as a movie they loved and would like to see something similar. By analyzing the knowledge graph, we can identify that ‘The Professional’ is an action, crime, and drama film with themes of revenge, vigilantism, and professional killers. The movie stars Burt Lancaster and Robert Duvall and is directed by Sidney Lumet. These elements suggest the user enjoys intense, character-driven narratives with a focus on crime and moral ambiguity. Overall preferences: Action, Crime, Drama, revenge, vigilantism, professional killer, intense character-driven narratives, moral ambiguity, Burt Lancaster, Robert Duvall, Sidney Lumet. Current interests: Action, Crime, Drama, revenge, professional killer, intense character-driven narratives.
@@
C.2. Case Study
@@
Table 5 presents a detailed visualization of a sample dialogue along with the corresponding preference summaries generated by different models. COMPASS demonstrates compelling effectiveness in capturing and reasoning about user preferences. It accurately identifies the movie genres and lists key cast members and the director of The Professional (1981). Moreover, COMPASS infers relevant themes that align with the movie’s content, such as professional killers, and intense action sequences. In contrast, other models show varying levels of accuracy and reasoning depth. (1) GPT-4o, without access to the KG, provides a broader interpretation. While it correctly identifies some genres, it also includes unsupported elements. This illustrates that while LLMs like GPT-4o possess some general world knowledge, they may lack the domain-specific information necessary for precise user preference summarization. (2) Llama-Summary, also without a knowledge graph, demonstrates significant inaccuracies. It incorrectly attributes the film to the French New Wave movement and lists incorrect personnel. This highlights the difficulties that smaller LLMs face in domain-specific tasks without specialized knowledge. (3) COMP w/o KG, while correctly identifying some genres and themes, introduces factual errors, such as misattributing the film to different actors and directors. This highlights the importance of integrating structured KGs, even when models are fine-tuned on domain-specific tasks. (4) COMP w/o GEP shows improved accuracy in genre identification but continues to struggle with factual details about the cast and crew. Although it makes plausible inferences, it still includes unsupported information. These comparisons emphasize the crucial roles that both KG integration and pre-training play in enabling COMPASS to generate accurate and contextually relevant user preference summaries.
Table 6. Instruction Templates for Graph Entity Captioning
Node Type Instruction Templates
Non-item nodes Please specify the node type (actor, director, genre, keywords, production company) and briefly provide associated details.
Specify the category of movie-related personnel or aspect and provide a concise summary.
Indicate the node type related to movies and offer a short overview of the associated key details.
What is the category of this node, and what are the main details related to it?
What is the type of movie-related entity, and what are some key details about it?
Item nodes Please summarize the details of the movie node, including its title, year, genre, keywords, director, and plot.
Provide a concise summary of the movie, including elements such as its title, release year, genre, key themes, director, and a brief plot overview.
Detail this movie’s specifics, including title, year, genre, keywords, director, and a short plot summary.
Present a summary of this film, noting its title, release year, genre, key themes, director, and plot.
Give a brief rundown of the movie, specifying the title, year it was made, genre, key terms, director, and plot.
Table 7. Instruction Template for User Preference Modeling and Recommendation
Instruction Template
As a movie recommender, analyze the user’s conversation history to make a movie recommendation. Break down the analysis into clear steps and return in json format:
1. reasoning: analyze the user’s conversation history and knowledge graph data to highlight their movie preferences.
2. overall preferences: A list of keywords summarizing the user’s overall preferences.
3. current interests: A list of keywords reflecting user’s current interests.
4. recommendation: a recommended movie title.
History: [Conversation history is inserted here]
Table 8. Instruction for Direct Recommendation
Instruction Template
As a movie recommender, please recommend a movie that aligns with the user’s interests discussed in the conversation history provided.
Conversation history: [Conversation history is inserted here]
Table 9. Caption Templates for Different Node Types
Node Type Caption Templates
Genre This node summarizes movies associated with the ¡genre¿ genre.
This node focuses on ¡genre¿ type movies.
Movies here are primarily categorized under the ¡genre¿ genre.
A collection of ¡genre¿ films defines this node.
This node represents a movie genre that is ¡genre¿.
Actor This node summarizes movies featuring actor ¡actor¿.
This node focuses on films starring ¡actor¿.
Movies with ¡actor¿ prominently featured make up this node.
This node collects major works of actor ¡actor¿.
This node highlights films that star ¡actor¿.
Director This node summarizes movies directed by ¡director¿.
This node focuses on films by director ¡director¿.
Films directed by ¡director¿ are centrally featured in this node.
This node is a compilation of ¡director¿’s directorial efforts.
A selection of ¡director¿’s films defines this node.
Company This node summarizes films produced by ¡company¿.
This node focuses on movies from the production company ¡company¿.
Productions by ¡company¿ are highlighted in this node.
This node represents the movie production company ¡company¿.
Films from ¡company¿ prominently make up this node.
Keyword This node summarizes movies associated with the keyword of ¡keyword¿.
This node focuses on movies themed around ¡keyword¿.
Films grouped in this node feature the theme: ¡keyword¿.
This node aggregates movies revolving around the keyword: ¡keyword¿.
This node indicates films related to the keyword: ¡keyword¿.
Table 10. Instruction Template for Ground Truth Generation
Instruction Template
As a personalized movie recommender, your task is to analyze the user’s conversation history and relevant movie data from the knowledge graph to discern a balanced mix of specific movie titles and broader thematic and stylistic movie preferences. Think step by step and your analysis should be nuanced and concise, using movie titles as entry points to uncover deeper interests in themes, genres, and styles.
# Analysis Steps
1. Review Conversation History: Extract mentions of movies and any expressed preferences in genres, keywords, actors, directors, or production companies from the dialogue.
2. Cross-Reference with Knowledge Graph: Correlate the extracted movie preferences from the conversation with the knowledge graph data, looking for patterns in genres, keywords, actors, directorial styles and etc.
3. Synthesize Insights: Combine the conversation and knowledge graph analysis to create a profile of the user’s movie preferences, highlighting both specific movies and general thematic interests.
# Input Example
## Conversation History
\$User:{}
\$Recommender: {}
## Knowledge Graph
(1). \${movie name}(\$year); Genres: \${Genre}; Keywords: \${Keywords}; Starring: \${Starring}; Director: \${Director}; Company: \${Company}; Plot: \${Plot}
# Expected Output
JSON Format Response
”reasoning”: ”A concise explanation reflecting the Analysis Steps, detailing how specific movie and broader thematic and stylistic preferences were identified and integrated.”,
”overall preferences”: ”A short list of keywords summarizing the user’s general thematic and stylistic movie preferences, including genres, notable keywords, preferred actors, directors, narrative styles, etc.”,
”current interest”: ”A concise list of keywords reflecting the user’s most recent and pertinent thematic and stylistic interests, guiding the next recommendations.”
Conversation History: [Conversation history is inserted here]
Knowledge Graph: [Mentioned Items Meta Information from Knowledge Graph is inserted here]

### v2

@@ V-B2 Comparison of Enhancement Methods
Figure 3: Comparison of recommendation performance on the INSPIRED dataset with different enhancers. Green percentages show improvements over baselines.
To evaluate COMPASS’s effectiveness, we compare it with Llama-Summary and GPT-4o as baseline enhancers, which generate user preference summaries without KG access. Figure 3 illustrates the performance of these methods across various CRSs on INSPIRED dataset. While all enhancement methods improve upon base models, COMPASS consistently outperforms both alternatives. Notably, COMPASS surpasses GPT-4o despite the latter being a much larger model, with improvements ranging from marginal gains on Llama3.1-8B to substantial improvements on KBRD and KGSF. This superior performance demonstrates the critical importance of integrating domain-specific knowledge into preference modeling. The substantial performance gap between COMPASS and Llama-Summary highlights the effectiveness of our framework in leveraging both structured knowledge and LLM capabilities. These results show COMPASS’s ability to generate more accurate and contextually relevant user preference summaries, leading to improved recommendation performance across different CRS architectures.
V-B3 Ablation study
The results in Table II show that COMPASS generally achieves the best performance across most metrics, highlighting its superior capability. The COM w/o REC variant shows the largest performance drop, emphasizing the crucial role of including recommended items in preference summaries to guide downstream models. Both COM w/o KG and COM w/o GEP variants demonstrate comparable performance declines, highlighting the crucial roles of KG integration for understanding item relationships and GEP for comprehending KG-augmented context. These findings show the importance of each COMPASS component, with their synergy driving the superior performance of the full model.
V-C Preference Generation Evaluation
To evaluate COMPASS’s ability to reason over user preferences and generate explainable summaries, we compare COMPASS with variants to assess key component effectiveness. Additionally, a case study with Llama-Summary and GPT-4o examples provides qualitative insights.
V-C1 Automatic Evaluation
TABLE II: Ablation study on recommendation task performance. The best-performing results are highlighted in bold.
@@ V-C1 Automatic Evaluation
TABLE III: Evaluation of generated user preference summaries. RP is reasoning proficiency and FC is factual consistency. The best results are highlighted in bold.
@@
TABLE IV: Case Study: Comparison of preference summaries generated by different models. Green indicates information supported by the KG, yellow indicates correct inferences not explicitly detailed in the KG but aligned with the content, and red indicates unsupported and unreasonable details.
@@
Table III presents a comparative analysis of COMPASS and its variants, COM w/o KG and COM w/o GEP, on ReDial and INSPIRED datasets. COMPASS consistently outperforms its variants across all metrics on both datasets. Lexical similarity metrics (ROUGE scores) demonstrate that COMPASS generates summaries that align closely with reference texts, while higher reasoning proficiency and factual consistency scores illustrate its superior reasoning quality and factual accuracy. The performance gap between COMPASS and COM w/o KG demonstrates the value of KG integration in enhancing reasoning over user preferences. Incorporating structured domain knowledge enables more accurate preference inference and explainable summary generation. Moreover, the comparative performance of COM w/o GEP and COMPASS indicates that, despite the presence of structured knowledge from the KG, there remains a modality gap between graph structures and natural language, posing challenges for LLMs to effectively reason with KG information. These results validate the COMPASS framework, demonstrating the synergistic effect of combining KGs with LLMs and confirming the critical role of our graph entity captioning pre-training in enabling effective cross-modal reasoning for explainable preference generation.
V-C2 LLM-Simulated User Evaluation
(a) ReDial: Alignment
(b) ReDial: Explainability
(c) INSPIRED: Alignment
(d) INSPIRED: Explainability
Figure 4: LLM-simulated user rankings for User Preference Alignment and Explainability across ReDial and INSPIRED datasets.
To assess user-perceived quality, we employ LLM-simulated user evaluations on explainability and user preference alignment. Figure 4 presents the results across ReDial and INSPIRED datasets. COMPASS is consistently preferred over COM w/o KG and COM w/o GEP in both metrics. On ReDial dataset, COMPASS is most preferred in approximately 75% of the cases for both alignment and explainability. Although the preference gap is smaller on INSPIRED, COMPASS maintains a clear advantage. Notably, COM w/o NP is consistently the least preferred, showing the importance of pre-training in generating meaningful preference summaries. These results validate COMPASS’s effectiveness in reasoning over user preferences and generating user-friendly, explainable summaries that enhance both fidelity and transparency in CRSs.
V-C3 Case Study
@@
We present a detailed case study in Table IV to demonstrate how different models reason over user preferences and generate explainable summaries from sample dialogue. COMPASS demonstrates compelling effectiveness in capturing and reasoning about user preferences. It accurately identifies the movie genres and lists key cast members and the director of The Professional (1981). Moreover, COMPASS infers relevant themes that align with the movie’s content, such as professional killers, and intense action sequences. In contrast, other models show varying levels of accuracy and reasoning depth. (1) GPT-4o, without KG augmentation, provides a broader interpretation but less precise reasoning. While it correctly identifies some genres, it includes unsupported elements, illustrating that general LLM knowledge alone is insufficient for accurate preference reasoning in domain-specific contexts. (2) Llama-Summary demonstrates significant inaccuracies, incorrectly attributing the film to the French New Wave movement and listing incorrect personnel. This highlights the limitations of smaller LLMs fin complex reasoning tasks without specialized knowledge. (3) COMP w/o GEP shows improved accuracy in genre reasoning but struggle with factual accuracy about cast and crew, making plausible but unsupported inferences. These comparisons emphasize the crucial roles that KG integration and proper pre-training play in enabling COMPASS to perform more accurate reasoning over user preferences, generating explainable summaries that enhance transparency and trust in CRSs.
VI Conclusion
In this paper, we introduce COMPASS, a novel framework that augments LLMs with KGs to reason over user preferences in CRSs. To address the modality gap between structured knowledge and natural language, we propose a graph entity captioning that transforms KG structures into LLM-compatible representations. Through knowledge-aware instruction tuning, COMPASS becomes proficient in performing cross-modal reasoning, generating interpretable user preference summaries. COMPASS has been extensively evaluated as a plug-and-play enhancement for various CRSs across benchmark datasets. The results demonstrate its effectiveness in significantly improving both recommendation performance and explainability of these base models. The adaptive integration mechanism of COMPASS allows for seamless enhancement of diverse CRS architectures without structural modifications, showcasing its versatility and potential for widespread adoption in the field of CRSs.